OpenStudy (anonymous):

Find derivative. f(x)=sin(x + x^1/2)

OpenStudy (anonymous):

(1/2)*cos(x+x^(1/2)) cos is the derivative of sin, and because of the chain rule, you must take the derivative of the inside (x+x^1/2) and multiply it to cos.

OpenStudy (anonymous):

oops! here's the right answer: [(1/2)x^(-1/2)]*cos(x+x^(1/2)) the above explanation still still correct.

OpenStudy (slaaibak):

\[{d \over{dx}} \sin(x + \sqrt{x}) = \cos(x + \sqrt{x}) * (1 + {1 \over {2\sqrt{x}}})\]

OpenStudy (slaaibak):

Janie you forgot to derive the x, hence you forgot the +1